OpenStudy (anonymous):

If tan(θ) = 5 7 , then sin(θ) = 5 and cos(θ) = 7. True or false?

OpenStudy (lgbasallote):

nice to know

OpenStudy (lgbasallote):

oh true or false..

OpenStudy (lgbasallote):

\[\tan \theta = \frac{\sin \theta}{\cos \theta}\] does that help?

OpenStudy (anonymous):

it's false.

OpenStudy (lgbasallote):

if your question says \[\text{If} \; \tan \theta = 5/7\] then it's true...

OpenStudy (anonymous):

not according to the answer key

OpenStudy (lgbasallote):

then the question is \[\tan \theta = 57\]???

OpenStudy (anonymous):

no. it's \[\tan \theta =5/7\]